Tohban Report 2012-03-07
From RHESSI Wiki
|Start Date:||6 Apr 2011|
|End Date:||13 Apr 2011|
|List all reports|
We have flares!! Activity is high this week with an X5 flare produced on 03/06/12. NOAA 11429 is a large beta-gamma-delta region with up to five significant penumbral spots. NOAA 11430 is an EFR thats growing just to the west of 11429 and may produce some sympathetic flaring events between the two regions. Hopefully activity won't die down until the region passes.
How many GOES flares occurred?
Flares above B, C, M, X class were 10 38 12 3
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 4 10 3 1
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 2 6 1 0
There were RHESSI flares/GOES flares 172 / 63 over the time range 29-Feb-12 07-Mar-12
We had major problems with the SSR this week. It reached a maximum of 99.6% during the X5. See below for explanation.
SSR is at 35% and continuing to empty as of the time of this message.
(Some of these were data dumps, some are not explained. )
TIME RANGE: 2012-03-06T00:00:00.000 -- 2012-03-07T00:00:00.000 NO GAPS IN APP_ID = 1 (VC1-SOH) NO GAPS IN APP_ID = 154 (VC1-PMTRAS) GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 6 GAP START TIME GAP END TIME GAP (SEC) 2012-03-06T03:30:00.000 -- 2012-03-06T03:45:00.000 900.00000 2012-03-06T04:00:00.000 -- 2012-03-06T05:05:00.000 3900.0000 2012-03-06T05:10:00.000 -- 2012-03-06T05:50:00.000 2400.0000 (2012-03-06T06:35:00.000 -- 2012-03-06T11:00:00.000 15900.000) (2012-03-06T11:25:00.000 -- 2012-03-06T14:50:00.000 12300.000) 2012-03-06T22:25:00.000 -- 2012-03-07T00:00:00.000 5700.0000
TIME RANGE: 2012-03-05T00:00:00.000 -- 2012-03-06T00:00:00.000 NO GAPS IN APP_ID = 1 (VC1-SOH) NO GAPS IN APP_ID = 154 (VC1-PMTRAS) GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 3 GAP START TIME GAP END TIME GAP (SEC) 2012-03-05T01:15:00.000 -- 2012-03-05T02:45:00.000 5400.0000 2012-03-05T02:55:00.000 -- 2012-03-05T03:15:00.000 1200.0000 (2012-03-05T04:00:00.000 -- 2012-03-05T19:25:00.000 55500.000)
NO GAPS IN APP_ID = 1 (VC1-SOH) NO GAPS IN APP_ID = 154 (VC1-PMTRAS) GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) (2012-03-04T11:20:00.000 -- 2012-03-05T00:00:00.000 45600.000)
DATA GAPS FOR 2012/03/02 TIME RANGE: 2012-03-02T00:00:00.000 -- 2012-03-03T00:00:00.000 NO GAPS IN APP_ID = 1 (VC1-SOH) NO GAPS IN APP_ID = 154 (VC1-PMTRAS) GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2012-03-02T15:35:00.000 -- 2012-03-02T15:40:00.000 300.00000
03/04/12 - 03/05/12 The SSR was noted to be at 96% on Sunday night and action was taken to change the decimation to active vigorous. This didn't have much effect to a days worth of Ago passes were added. The fast rates were accidentally set to force on instead of auto on when it was noted they weren't turned back on after the anneal.
It was established that the attenuators were not moving and so the thick attenuator was forced in place. Following an ADP reset, the attenuators were found to be working automatically again. However, just to be on the safe side, the two attenuators were forced in to try to empty out the SSR which was rising faster than we could empty (despite all the extra passes). Finally we decided to dump data to make space in the SSR which was back at 90%.
It was thought that the increased data rate was as a result of increased number of counts in the unsegmented G2 and G4. The thresholds of G2 and G4 were increased to ff (maximum) in order to inhibit as many counts as possible and G2 and G4 were set to not take events during the high magnetic latitude zones.
Finally, Jeremy realized that the data collection rate was increasing during times of eclipse. Mark hypothesized that since livetime = 0 in eclipse, the fast rates were being triggered and increasing our data collection. The fast rates were forced off and the SSR has been falling since.
Since the SSR filling has been attributed to the fast rates, the following steps will be followed to undo the precautions put in place over the past few days:
- Make sure the fast rates remain disabled.
- Return thick attenuator to automatic.
- Reduce the thresholds of G2 and G4 to what they were.
- Depending on the SXR flux background level, return thin shutter motions to auto.
- Reinstate G2 and G4 events in high magnetic latitude zones.